# Seeing through Satellite Images at Street Views

### v1

@@ 3.3 Training-Time Satellite-View Generation
To enhance the appearance alignment between the learned 3D representation and input $I_{\text{sat}}$, we render the satellite view during training to provide additional supervision. Since existing datasets [11, 5, 35] did not provide satellite camera intrinsic and extrinsic parameters, we approximate the satellite camera as an orthographic camera oriented vertically downward toward the ground. With the camera approximation, we can compute the ray origins and ray directions corresponding to each pixel for the satellite view rendering.
Unlike street-view image generation, the sky generation module and SR module are not required when generating satellite-view images. To deal with the absence of street-view grounded illumination, when synthesizing satellite-view images, we substitute the illumination style vector $\mathbf{w}_{\text{ill}}$ in Equ. 4 with a null-style vector $\mathbf{w}_{0}=\mathbf{0}$, as the tri-plane feature map $F_{\text{tri}}$ is directly derived from input satellite imagery. This zero-vector serves as a fixed placeholder to disable unnecessary style modulation in the generation pipeline. Finally, given satellite camera parameters, the model can generate satellite-view image feature $\hat{I}_{F\text{sat}}$, image color $\hat{I}_{C\text{sat}}$, depth $\hat{D}_{\text{sat}}$ by volume rendering.
@@ 3.4 Learning Objectives
The training data consists of paired satellite and street-view images, we also provide extra binary pseudo sky masks $M_{\text{sky}}$ for each street-view images by an off-the-shelf sky segmentation model [30]. In this section, we introduce the learning objectives, including reconstruction, opacity, and discriminator losses.
@@ 3.4.1 Non-sky Opacity Loss
In Sat2Density++, we aim to learn the 3D representation of the satellite scene, but street-level scenes introduce elements such as the sky, which are not accounted for in satellite images. This challenges the model’s ability to recognize and accurately model the geometric structure of scenes captured in satellite imagery. We found the pseudo sky masks provide a strong inductive basis to regularize the density field in a simple way, thus using a binary cross-entropy (BCE) loss for the supervision:
@@
where $M_{\text{sky}}$ is the ground-truth sky mask, 1 indicates the sky region and 0 indicates the ground region.
@@ 3.4.2 Reconstruction Loss
where $\lambda_{\text{sat}}$, $\lambda_{\text{str}}$, and $\lambda_{\text{sky}}$ balance three terms. Besides, for save consumption, when synthesizing satellite view images, we only render at $1/4$ resolution of input satellite image size by random crop during training.
@@ 3.4.3 Discriminator Loss
The reconstruction loss alone fails to synthesize detailed results from novel viewpoints. Therefore, we use an adversarial loss [37] to enforce renderings to look realistic from rendered viewpoints. Specifically, we have two discriminators $D_{\rm str}$ and $D_{sat}$ for street-view images and satellite view images, respectively.
@@
$\mathcal{L}_{\text{GAN}}$ $=\lambda_{\text{\rm D\_str}}\mathcal{L}_{\rm D\_str}(\hat{I}_{C\text{str}}^{+},\hat{I}_{C\text{str}})+\lambda_{\rm D\_sat}\mathcal{L}_{D_{\text{sat}}}(\hat{I}_{C\text{sat}}),$ (18)
where $\lambda_{\text{pano}}$ and $\lambda_{\text{sat}}$ are hyperparameters for loss weighting, $\hat{I}_{C\text{str}}$ is the first three layers of blended street-view panorama raw feature $\hat{I}_{F\text{str}}$. The non-saturating GAN loss function [37] with R1 regularization [38] is used for $\mathcal{L}_{\text{GAN}}$, following the training scheme in StyleGAN2 [31].
@@ 3.4.4 Total Loss
Figure 3: Three video results generated by our method and Sat2Density[1] on the VIGOR dataset. The full videos can be seen on the project page.
Figure 4: Three video results generated by our method and Sat2Density [1] on the CVACT dataset. The full videos can be seen on the project page.
@@ 4.2 Implementation Details
Our code is built on Carver [40], an efficient PyTorch-based library for training 3D-aware generative models. For hyperparameters, we empirically assign the values of $\lambda_{\rm D\_str}$, $\lambda_{\rm D\_sat}$, $\lambda_{\text{sat}}$, $\lambda_{\text{str}}$, $\lambda_{\text{sky}}$, and $\lambda_{\text{opa}}$ to 1.0, 1.0, 30.0, 10.0, 10.0, and 25.0, respectively. We use 8 NVIDIA A100 GPUs for training, with a batch size of 4 per GPU. For the CVACT and CVUSA datasets, the training involves 180,000 iterations and takes about 30 hours. In contrast, training on the VIGOR dataset involves 360,000 iterations over an extended duration of approximately 60 hours.
@@ 4.3 Evaluation Protocols and Metrics
The evaluation metrics used in our experiments comprehensively measure the quality of the generated street-view images and videos, which include FID and KID for realism evaluation and DINOv2 similarity for the high-level consistency, as well as the conventional used SSIM [41], PSNR [42] and LPIPS [36]. For the evaluation of generated videos, due to the absence of video groundtruth, user studies on the CVACT and VIGOR datasets are used to measure the quality, consistency, and faithfulness of videos.
@@
SSIM, PSNR and LPIPS. These three metrics follow the previous works [5, 1] to measure the quality of generated images. SSIM assesses the structural similarity, PSNR gauges the pixel-wise distance, and LPIPS [36] computes the high-level perceptual similarity between two images. For the computation of LPIPS, we utilize the pre-trained AlexNet [47] and SqueezeNet [48] as the feature backbones, referred to as $P_{\text{alex}}$ and $P_{\text{squ}}$, respectively.
User Studies for Generated Video. Existing datasets consist solely of image collections and lack video data, so we quantitatively compare video results through a user study involving 31 participants and comprising 40 sets of videos. Each set is based on an input satellite image and camera trajectory video and includes videos generated by both our method and Sat2Density. 20 video sets are derived from the CVACT dataset, while the remaining from the VIGOR dataset. Specifically, we assess the faithfulness of the generated videos to the input satellite images (Faithfulness), the consistency of the videos across multiple views (Consistency), and the overall visual quality exhibited by the generated content (Quality).
@@ 4.4 Main Comparisons
We compare Sat2Density++ with both image generation and video generation methods. For image generation methods, we evaluate against Pix2Pix [6], XFork [17], and Shi et al. [5]. For video generation methods, we compare with our preliminary work Sat2Density [1], which synthesizes street-view images from arbitrary camera positions. Note that on the VIGOR dataset, the positions of street-view images do not always located at the center of the input satellite images. Therefore, we limit our comparisons with methods [6, 17, 5], which can only synthesize single images at the center of the input satellite images, to the CVACT and CVUSA datasets.
TABLE I: Quantitative comparison on image level metric with existing algorithms on the CVUSA, CVACT, VIGOR, and VIGOR-ood test set. For Sat2Density and Ours, ‘*’ indicates that for each satellite image input, we randomly select an illumination feature from the training set as the illumination input for rendering street views, which provides a fair comparison and ‘†’ means we input ground-truth illumination feature, shows the effect of illumination feature on the metric.
@@
4.4.1 Street-View Panorama Video Generation
@@
To the best of our knowledge, beyond Sat2Density and our proposed Sat2Density++, only Sat2Vid[9] and Sat2Scene[10] aim to generate videos from satellite imagery. However, both rely on building height maps and focus primarily on rendering facades and road surfaces, limiting scene diversity. Due to the unavailability of their full datasets and code, we are unable to conduct comparisons with them.
@@
Overall, unlike prior work that focuses narrowly on buildings, our method generates a diverse range of scene elements—including trees, green belts, and landmarks—resulting in videos that are more faithful to satellite inputs and visually realistic. Moreover, Sat2Density++ does not require additional supervision such as height maps or metric depths, making it more versatile and broadly applicable.
4.4.2 Street-View Image Synthesis
@@
The quantitative image-level results in Tab. I demonstrate the superior performance of our method. Across all four test sets—even when using randomly sampled illumination inputs from the training set (denoted by ‘*’)—our approach consistently outperforms existing methods in perceptual, semantic, and realism metrics. Notably, Sat2Density++ achieves greater gains over Sat2Density on the complex urban test sets (VIGOR-ood and VIGOR) than on the simpler suburban ones, highlighting its improved capability in modeling complex scenes. These improvements stem from our more effective 3D representation learning design.
@@
While Sat2Density++ shows slightly lower PSNR and SSIM scores than Sat2Density on CVUSA and CVACT, and also underperforms Shi et al. [5] on CVUSA, this does not indicate true performance drops. Instead, it reflects the limitations of low-level metrics, which are overly sensitive to appearance variations in homogeneous regions—common in suburban panoramas. This is further evidenced by the comparison using real versus random illumination inputs (denoted by ‘†’ in Tab. I). With real illumination features, perceptual, PSNR, and SSIM scores improve significantly, while realism metrics such as FID and KID remain stable. This demonstrates the sensitivity of low-level metrics to appearance changes, even when scene semantics are preserved. Realism metrics, based on distributional similarity, offer a more robust measure of perceptual quality. Together, these findings confirm that Sat2Density++ effectively captures and utilizes illumination information, enabling visually consistent and illumination-controllable image synthesis, as illustrated in Fig.9 and Fig.8.
@@
Figure 6: Ablation study for the proposed strategies on the VIGOR dataset. The first column is the input satellite image, the cross point in the satellite image is the input camera position, the masked sky image for illumination input, and ground truth street-view image. The red boxes show the failure regions. The video ablation can be seen on the project page.
TABLE II: Ablation study for illumination input on VIGOR-ood test set for satellite-view and street-view generation. In this experiment, the illumination input for satellite and street views are set to random that sample a random vector from Gaussian distribution, zero-style that is a zero vector, and valid that use the ground-truth illumination input during training.
@@
Zero-Style Random 0.439 64.98 .060
Zero-Style Valid 0.465 40.85 .035
TABLE III: Ablation results on VIGOR-ood test set. ‘w/o $\mathcal{L}_{\text{opa}}$’ means training without non-sky opacity loss, ‘w/o illu.’ means no illumination input $f_{\text{ill}}$, and replace the $f_{\text{ill}}$ and $\mathbf{w}_{0}$ to a random noise $\mathbf{z}$. ‘w/o $\mathcal{L}_{\text{sky}}$’ means removing the sky reconstruction loss. ‘w/o sat. view loss’ means we remove the sat. discriminator and sat. view reconstruction loss.
@@ 4.5 Ablation Study
In this section, we will discuss the contributions of key components within Sat2Density++. The ablation study was conducted on the VIGOR-ood test set. We present the quantitative ablation in Tab. II, Tab. III, Tab. IV and qualitative ablation in Fig. 6 and Fig. 7, Fig.8.
4.5.1 Illumination Modeling
The illumination input, illumination-adaptive tri-plane decoder, and illumination-adaptive 2D sky generation module work together to mitigate the effects of view-specific lighting in street-view images during training, while also enabling controllable illumination synthesis at inference. Tab. II summarizes different design choices for modeling illumination in the satellite and street-view branches.
As shown in Tab. II, the poorest performance occurs when both satellite and street views are generated using random noise, completely ignoring the inherent illumination differences between the two views. The corresponding visual results are shown in the “w/o illu.” column of Fig. 6. Compared to the full model, the predicted geometry from the satellite view shows larger deviations along object boundaries and more angular artifacts in depth, as highlighted in the red box in the first row. This demonstrates that failing to account for differing imaging conditions leads to degraded 3D representations and lower-quality video generation. A slight improvement is observed when using the zero-style for satellite-view generation, while still using random noise for the street view. In this case, the model becomes aware of cross-view appearance differences, but lacks proper compensation through illumination-aware modeling in the street view.
Our full model, which uses zero-style for satellite-view generation and the valid sky histogram for the street-view illumination input, achieves the best performance. This configuration fully leverages illumination cues from both views, leading to more accurate geometry, better visual quality, and controllable street-view synthesis.
Figure 7: Ablation study on satellite view loss.“w/o sat. view loss” denotes training without $\mathcal{L}_{\text{sat}}$ and $\mathcal{L}_{D_{\text{sat}}}$.

### v2

@@ 3.3 Training-Time Satellite-View Generation
To enhance the appearance alignment between the learned 3D representation and input $I_{\text{sat}}$, we render the satellite view during training to provide additional supervision. Since existing datasets [67, 46, 72] did not provide satellite camera intrinsics and extrinsics, we approximate the satellite camera as an orthographic camera oriented vertically downward toward the ground. With the camera approximation, we can compute the ray origins and ray directions corresponding to each pixel for the satellite view rendering.
Unlike street-view image generation, the sky generation module and SR module are not required when generating satellite-view images. To deal with the absence of street-view grounded illumination, when synthesizing satellite-view images, we substitute the illumination style vector $\mathbf{w}_{\text{ill}}$ in Eq. (4) with a null-style vector $\mathbf{w}_{0}=\mathbf{0}$, as the tri-plane feature map $F_{\text{tri}}$ is directly derived from input satellite imagery. This null-style vector serves as a fixed placeholder to disable unnecessary style modulation in the generation pipeline. Finally, given satellite camera parameters, the model can generate satellite-view image feature $\hat{I}_{F\text{sat}}$, image color $\hat{I}_{C\text{sat}}$, depth $\hat{D}_{\text{sat}}$ by volume rendering.
@@ 3.4 Learning Objectives
The training data consists of paired satellite and street-view images. We also provide extra binary pseudo sky masks $M_{\text{sky}}$ for each street-view image by an off-the-shelf sky segmentation model [69]. In this section, we introduce the learning objectives, including reconstruction, opacity, and discriminator losses.
@@ 3.4.1 Non-sky Opacity Loss
In Sat2Density++, we aim to learn the 3D representation of the satellite scene, but street-level scenes introduce elements such as the sky, which are not accounted for in satellite images. This challenges the model’s ability to recognize and accurately model the geometric structure of scenes captured in satellite imagery. We found that the pseudo sky masks provide a strong inductive basis to regularize the density field in a simple way, thus using a binary cross-entropy (BCE) loss for the supervision:
@@
where $M_{\text{sky}}$ is the ground-truth sky mask, 1 indicates the sky region, and 0 indicates the ground region.
@@ 3.4.2 Reconstruction Loss
where $\lambda_{\text{sat}}$, $\lambda_{\text{str}}$, and $\lambda_{\text{sky}}$ balance three terms. Besides, to save consumption, when synthesizing satellite view images, we render at one-quarter resolution with random crops during training.
@@ 3.4.3 Discriminator Loss
The reconstruction loss alone fails to synthesize detailed results from novel viewpoints. Therefore, we use an adversarial loss [9] to enforce renderings to look realistic from rendered viewpoints. Specifically, we have two discriminators $D_{\rm str}$ and $D_{sat}$ for street-view images and satellite-view images, respectively.
@@
$\mathcal{L}_{\text{GAN}}$ $=\lambda_{\rm D_{\text{str}}}\mathcal{L}_{\rm D_{\text{str}}}(\hat{I}_{C\text{str}}^{+},\hat{I}_{C\text{str}})+\lambda_{\rm D_{\text{sat}}}\mathcal{L}_{\rm D_{\text{sat}}}(\hat{I}_{C\text{sat}}),$ (18)
where $\lambda_{\rm D_{\text{str}}}$ and $\lambda_{\rm D_{\text{sat}}}$ are hyperparameters for loss weighting, $\hat{I}_{C\text{str}}$ is the first three layers of blended street-view panorama raw feature $\hat{I}_{F\text{str}}$. The non-saturating GAN loss function [9] with R1 regularization [36] is used for $\mathcal{L}_{\text{GAN}}$, following the training scheme in StyleGAN2 [17].
@@ 3.4.4 Total Loss
Figure 3: Three video results generated by our method and Sat2Density[41] on the VIGOR dataset. For each example, the left panel shows the satellite image with the camera trajectory, which goes from the red point to the blue point. Along this trajectory, we select five camera positions (highlighted points) to illustrate multi-view continuity. The five groups of panoramas on the right correspond to these five camera positions in order. In each group, the upper row shows the results of Sat2Density, while the lower row shows the corresponding results of our method at the same camera positions, ensuring a fair one-to-one comparison. The full videos can be seen on the project page.
Figure 4: Three video results generated by our method and Sat2Density [41] on the CVACT dataset. For each example, the left panel shows the satellite image with the camera trajectory, which goes from the red point to the blue point. Along this trajectory, we select five camera positions (highlighted points) to demonstrate the multi-view continuity around roundabouts or road intersections. The five groups of panoramas on the right correspond to these five camera positions in order. In each group, the upper row shows the results of Sat2Density, while the lower row shows the corresponding results of our method at the same camera positions, ensuring a fair one-to-one comparison. The full videos, including all sampled positions, can be found on the project page.
@@ 4.2 Implementation Details
Our code is built on Carver [55], an efficient PyTorch-based library for training 3D-aware generative models. For hyperparameters, we empirically assign the values of $\lambda_{\rm D_{\text{str}}}$, $\lambda_{\rm D_{\text{sat}}}$, $\lambda_{\text{sat}}$, $\lambda_{\text{str}}$, $\lambda_{\text{sky}}$, and $\lambda_{\text{opa}}$ to 1.0, 1.0, 30.0, 10.0, 10.0, and 25.0, respectively. We use 8 NVIDIA A100 GPUs for training, with a batch size of 4 per GPU. For the CVACT and CVUSA datasets, the training involves 180,000 iterations and takes about 30 hours. In contrast, training on the VIGOR dataset involves 360,000 iterations over an extended duration of approximately 60 hours.
@@ 4.3 Evaluation Protocols and Metrics
The evaluation metrics used in our experiments comprehensively measure the quality of the generated street-view images and videos, which include FID and KID for realism evaluation and DINOv2 similarity for the high-level consistency, as well as the conventionally used SSIM [57], PSNR [56], and LPIPS [71]. For the evaluation of generated videos, due to the absence of video ground truth, user studies on the CVACT and VIGOR datasets are used to measure the quality, consistency, and faithfulness of videos.
@@
SSIM, PSNR, and LPIPS. These three metrics follow the previous works [46, 41] to measure the quality of generated images. SSIM assesses the structural similarity, PSNR gauges the pixel-wise distance, and LPIPS [71] computes the high-level perceptual similarity between two images. For the computation of LPIPS, we utilize the pre-trained AlexNet [19] and SqueezeNet [14] as the feature backbones, referred to as $P_{\text{alex}}$ and $P_{\text{squ}}$, respectively.
User Studies for Generated Video. Existing datasets consist solely of image collections and lack video data, so we quantitatively compare video results through a user study involving 31 participants and comprising 40 sets of videos. Each set is based on an input satellite image and camera trajectory video and includes videos generated by both our method and Sat2Density. 20 video sets are derived from the CVACT dataset, while the remaining are from the VIGOR dataset. Specifically, we assess the faithfulness of the generated videos to the input satellite images (Faithfulness), the consistency of the videos across multiple views (Consistency), and the overall visual quality exhibited by the generated content (Quality).
@@ 4.4 Main Comparisons
We compare Sat2Density++ with both image generation and video generation methods. For image generation methods, we evaluate against Pix2Pix [15], XFork [43], and Shi et al. [46]. For video generation methods, we compare with our preliminary work Sat2Density [41], which synthesizes street-view images from arbitrary camera positions. Note that on the VIGOR dataset, the positions of street-view images are not always centered in the input satellite patch. Therefore, we limit our comparisons with methods [15, 43, 46], which can only synthesize single images at the center of the input satellite images, to the CVACT and CVUSA datasets.
TABLE I: Quantitative comparison on image level metric on the CVUSA, CVACT, VIGOR, and VIGOR-OOD test set. For Sat2Density and Ours, ‘*’ indicates that for each satellite image input, we randomly select a sky illumination feature from the training set as the illumination input for rendering street views, which provides a fair comparison, and ‘†’ means we input ground-truth sky illumination feature, shows the effect of illumination feature on the metric.
@@
4.4.1 Comparison on Video Results
@@
To the best of our knowledge, beyond Sat2Density and our proposed Sat2Density++, only Sat2Vid [30] and Sat2Scene [29] aim to generate videos from a satellite image. However, both rely on building height maps and focus primarily on rendering facades and road surfaces, limiting scene diversity. Due to the unavailability of their full datasets and code, we are unable to conduct comparisons.
@@
Overall, unlike prior work that focuses narrowly on buildings, our method generates a diverse range of scene elements, including trees, green belts, and landmarks, resulting in videos that are more faithful to satellite inputs and visually realistic. Moreover, Sat2Density++ does not require additional supervision such as height maps or metric depths, making it more versatile and broadly applicable.
Figure 6: Sky generation from sky-illumination inputs for (a) in-domain and (b) OOD scenarios. In each pair, the left shows a real sky image, while the right presents a sky synthesized from the extracted sky-region histogram. In-domain cases may display higher resemblance due to near-time captures relative to the training data, while OOD cases consistently maintain plausible tone, lighting, and cloud distribution, demonstrating robust generalization of the sky generation.
4.4.2 Comparison on Image Results
@@
The quantitative image-level results in Tab. I demonstrate the superior performance of our method. Across all four test sets, even when using randomly sampled illumination inputs from the training set (denoted by ‘*’), our approach consistently outperforms existing methods in perceptual, semantic, and realism metrics. Notably, Sat2Density++ achieves greater gains over Sat2Density on the complex urban test sets (VIGOR-OOD and VIGOR) than on the simpler suburban ones, highlighting its improved capability in modeling complex scenes. These improvements stem from our more effective 3D representation learning design.
@@
While Sat2Density++ shows slightly lower PSNR and SSIM scores than Sat2Density on CVUSA and CVACT, and also underperforms Shi et al. [46] on CVUSA, this does not indicate true performance drops. Instead, it reflects the limitations of low-level metrics, whose sensitivity to appearance variations in homogeneous regions becomes especially problematic in suburban panoramas. This is further evidenced by the comparison using real versus random illumination inputs (denoted by ‘†’ in Tab. I). With real illumination features, perceptual, PSNR, and SSIM scores improve significantly, while realism metrics such as FID and KID remain stable. This demonstrates the sensitivity of low-level metrics to appearance changes, even when scene semantics are preserved. Realism metrics, based on distributional similarity, offer a more robust measure of perceptual quality. Together, these findings confirm that Sat2Density++ effectively captures and utilizes illumination information, enabling visually consistent and illumination-controllable image synthesis, as illustrated in Fig. 12 and Fig. 9.
@@
Figure 7: Ablation study for the proposed strategies on the VIGOR dataset. The first column is the input satellite image, the cross point in the satellite image is the input camera position, the masked sky image for sky illumination input, and the ground truth street-view image. The red boxes show the failure regions. The video ablation can be seen on the project page.
TABLE II: Ablation of illumination inputs on VIGOR-OOD test set for satellite- and street-view generation. Each row sets the illumination code for the two branches (columns “Satellite” and “Street View”): random draws a vector $z\!\sim\!\mathcal{N}(0,I)$, null-style uses the null-style vector $\mathbf{w}_{0}$, and valid supplies the ground-truth illumination when training. Bold denotes the best configuration (Satellite=null-style, Street=valid).
@@
Null-Style Random 0.439 64.98 .060
Null-Style Valid 0.465 40.85 .035
TABLE III: Ablation results on VIGOR-ood test set. ‘w/o $\mathcal{L}_{\text{opa}}$’ means training without non-sky opacity loss, ‘w/o illu.’ means no illumination input $f_{\text{ill}}$, and replace the $f_{\text{ill}}$ and $\mathbf{w}_{0}$ with a random noise $\mathbf{z}$. ‘w/o $\mathcal{L}_{\text{sky}}$’ means removing the sky reconstruction loss. ‘w/o sat. view loss’ means removing the sat. discriminator and sat. view reconstruction loss. ‘w/o sky branch’ means we remove the sky generation branch, use the tri-plane NeRF to render a full street-view image. ‘w. vanilla NeRF’ means we replace the illumination adaptive tri-plane decoder with a vanilla tri-plane decoder.
@@
4.5 Generalizability of sky generation
We included several visualizations of sky generation from sky illumination input on both in-domain and out-of-domain sets, as shown in Fig. 6. In the VIGOR in-domain test set, some of the training and test data were captured on the same day under nearly identical lighting conditions. Consequently, it is possible to find cases where the generated sky image closely resembles the original sky image corresponding to the sky illumination input, as the training set might have encountered the same sky illumination input. In out-of-domain testing, however, such cases do not occur. The color histograms of sky regions not seen during training can still produce well-rendered sky areas. We observe a high similarity in sky lighting between the generated and real images, demonstrating the robust generalization capability of our sky illumination input.
@@ 4.6 Ablation Study
In this section, we will discuss the contributions of key components within Sat2Density++. The ablation study was conducted on the VIGOR-OOD test set. We present the quantitative ablation in Tab. II, Tab. III, Tab. IV and qualitative ablation in Fig. 7 and Fig. 8, Fig. 9. Unless otherwise stated, all quantitative comparisons in the ablation study are conducted under the same setting as ’†’ in Tab. I.
4.6.1 Sky Illumination Modeling
The sky illumination input, illumination-adaptive tri-plane decoder, and illumination-adaptive 2D sky generation module work together to mitigate the effects of view-specific lighting in street-view images during training, while also enabling controllable illumination synthesis at inference. Tab. II summarizes different design choices for modeling illumination in the satellite and street-view branches.
As shown in Tab. II, the poorest performance occurs when both satellite and street views are generated using random noise, completely ignoring the inherent illumination differences between the two views. The corresponding visual results are shown in the “w/o illu.” column of Fig. 7. Compared to the full model, the predicted geometry from the satellite view shows larger deviations along object boundaries and more angular artifacts in depth, as highlighted in the red box in the first row. This demonstrates that failing to account for differing imaging conditions leads to degraded 3D representations and lower-quality video generation. A slight improvement is observed when using the null-style vector for satellite-view generation, while still using random noise for the street view. In this case, the model becomes aware of cross-view appearance differences, but lacks proper compensation through illumination-aware modeling in the street view.
Our full model, which uses a null-style vector for satellite-view generation and the valid sky histogram for the street-view illumination input, achieves the best performance. This configuration fully leverages illumination cues from both views, leading to more accurate geometry, better visual quality, and controllable street-view synthesis.
Figure 8: Ablation study on satellite-view loss.“w/o sat. view loss” denotes training without $\mathcal{L}_{\text{sat}}$ and $\mathcal{L}_{D_{\text{sat}}}$.